OpenStudy (anonymous):

answers to math problems Montel can't keep up with the demand for the latest best-selling book. He had to order more copies twice this week. On Monday he placed an order for 96 books. On Wednesday he ordered 264 more. All the books arrived on Friday. There were 24 books in each box. How many boxes of books did Montel receive?

OpenStudy (anonymous):

Total books ordered = 456 So, no. of boxes received = 456/24 = 19 boxes

OpenStudy (anonymous):

medals???

OpenStudy (anonymous):

15 boxes

OpenStudy (anonymous):

how did you get 15?

OpenStudy (anonymous):

Total ordered books = 360

OpenStudy (anonymous):

But it is mentioned 264 MORE books right?

OpenStudy (anonymous):

So total no. of ordered books = 96 + 96 + 264

OpenStudy (anonymous):

But it was not given before that he was going to buy 96 books on Wednesday.

OpenStudy (hoblos):

i go with abdul shabeer 15

OpenStudy (anonymous):

montel?